OpenStudy (anonymous):

HELP!! MEDAL WILL BE GIVEN!!! Use a suitable rotation of axes to find the equation for the graph in an x'y'-plane and sketch the graph, labeling the vertices. Show work on how you got the answer.

OpenStudy (anonymous):

\[x ^{2}-2xy+y ^{2}-2\sqrt{2}x-2\sqrt{2}y=0\]

OpenStudy (anonymous):

@Hero @jdoe0001 @Directrix

OpenStudy (anonymous):

Anybody?

OpenStudy (jdoe0001):

hmm dunno what's meant by "rotation of the axes", I can tell is an ellipse pretty much

OpenStudy (anonymous):

Well I learned was that you have to find an angle form using \[\cot 2\]

OpenStudy (jdoe0001):

hmmm.... don't think I've covered that

OpenStudy (anonymous):

Well our professor told us to start with\[\frac{ a-c }{ b }\]then that equals to cot 2 theta somehow

OpenStudy (jdoe0001):

hmmm dunno that one :(

OpenStudy (anonymous):

Its ok I already missed the homework deadline